# DeNOTS: Stable Deep Neural ODEs for Time Series

### v4

@@ C.1 Datasets
  • The PhysioNet Sepsis  [^6] https://physionet.org/content/challenge-2019/1.0.0/ dataset Reyna et al. (2020), released under the CC-BY 4.0 license. Most of the values are missing, with only 10% present. Sequences are relatively short, ranging from 8 to 70 elements, with a median length of 38. The target variable indicates whether the patient developed sepsis, meaning it’s binary and highly unbalanced, so we measure AUROC. This dataset is also used by the original NCDE paper Kidger et al. (2020), making it a valuable addition to our benchmark. Following the NCDE paper, we only consider the first 72 hours of a patient’s stay.
  • Two datasets from the Time Series Classification archive  [^7] https://www.timeseriesclassification.com/ Middlehurst et al. (2024) (with no explicit license specified): the U-Wave Gesture Library Liu et al. (2009) and the Insect Sound dataset [^8] https://www.timeseriesclassification.com/description.php?Dataset=InsectSound. These datasets were processed identically: the timestamps were taken to be evenly spaced between 0 and $1$, and 30% of each sequence was replaced with NaNs. They both have a multiclass label, so we use the Accuracy metric.
  • A synthetic Pendulum dataset, with the dampening coefficient as the target Osin et al. (2024) (with no explicit license specified) — regression task. Sequences are irregular, containing 10% missing values, with lengths varying from approximately 200 to 400 elements, the median length being 315.
@@ C.2 Backbones
The recurrent networks are all single-layer and one-directional, with 32 hidden channels. For GRU Chung et al. (2014) we use the standard PyTorch implementation:
@@
We use the Pytorch implementation of the Transformer Encoder layer. The two versions differ in their positional embeddings: RoFormer uses rotary embeddings Su et al. (2024) (provided by TorchTune), and TempFormer uses temporal sine-based embeddings, as proposed by the Transformer Hawkes paper Zuo et al. (2020).
@@
The Mamba model we use is Mamba2 Gu & Dao (2023) from the Mamba SSM library  [^9] https://github.com/state-spaces/mamba, which we employ without significant modifications.
@@
To reproduce the results as closely as possible, we use TorchCDE  [^10] https://github.com/patrick-kidger/torchcde for the Neural CDE method Kidger et al. (2020). Specifically, we implemented the example for irregular sequence classification from their repository. For Neural RDE Morrill et al. (2021b), we pre-process the data using the original Signatory package, provided by the authors  [^11] https://github.com/patrick-kidger/signatory, with depth set to 2.
@@
As the model for GRU-ODE De Brouwer et al. (2019), we simply take a Neural CDE with the Sync-NF vector field, without time scaling.
@@
DeNOTS is slower than other non-ODE methods; however, we argue this is an implementation issue. Neural ODEs received much less attention than State-Space Models, Transformers, or RNNs. We use the TorchODE library Lienen & Günnemann (2022). Although it is faster than the original implementation Chen et al. (2018), it still requires work, being built almost single-handedly by its main contributor.
@@ C.3 Pipeline
We provide our repository with automated scripts to download and preprocess all datasets, as well as to train and evaluate all the considered models, using popular frameworks such as Pytorch Ansel & team (2024) (Caffe2 license), Pytorch Lightning Falcon & The PyTorch Lightning team (2019) (Apache License 2.0) and Hydra Yadan (2019) (MIT License) to make the process familiar to most AI researchers. The YAML configs, containing all the hyperparameters, are also provided in the repository to facilitate reproducibility. The documentation clearly explains the steps necessary to reproduce all considered experiments.
@@ C.3.1 Training
All training uses the Adam method Kingma (2014), with the learning rate fixed to $10^{-3}$ and other parameters left default. The whole model, consisting of the head and the backbone, is trained end-to-end. The head consists of a linear layer and an activation function. We do not set an upper bound for epochs, stopping only when the validation metric stops improving. The choice of head’s activation, loss, and specific metrics depends on the considered task, as outlined in Table 12.
@@ C.3.2 Embeddings
Before passing $\{\mathbf{x}_{k}\}_{k}$ to the backbone, we apply Batch Normalization Ioffe & Szegedy (2015). For a fairer baseline comparison, time intervals $t_{k}-t_{k-1}$ are also included in the embeddings in the same way as other numerical features.
@@
C.4 Expressivity discussion
Here, we explain why measuring expressivity with the downstream metric is a valid approach. Theoretically, the total error can be decomposed into three terms Gühring et al. (2020):
@@ C.4 Expressivity discussion
Next, it is the consensus that the more expressive models suffer more from overfitting Hawkins (2004), i.e. generalize worse. We verify that our models adhere to this by performing the following experiment. We construct a smaller version of the Pendulum dataset; its training set is 1/32 of the original one, and compare the DeNOTS model with various values of $D$ on this benchmark. The results are provided in Table 13. The overfitting effect is observed very clearly: the ”shallow“ $D=1$ model achieves higher metrics on the test set than $D=10$, and the ranking is reversed on the train set.
@@ C.5 NFE-Correlation experiments
C.6 Negative Feedback Strength
@@ C.7.2 Sine-2 Dataset
C.8 Full attack results
@@
Additionally, we test how the value of $D$ affects the robustness of our models. This is presented in Figures $25$. Increasing depth lowers robustness, which is in line with the consensus. Prior work indicates that the more accurate models with higher metrics often perform worse under adversarial attacks than those with lower metrics Su et al. (2018). Increased accuracy often requires increased expressiveness and sensitivity, which directly impacts robustness. Consequently, this demonstration supports our claim that larger values of $D$ correspond to more expressive models.

### v5

@@ C.1 Datasets
  • The PhysioNet Sepsis  [^4] https://physionet.org/content/challenge-2019/1.0.0/ dataset reyna2020early, released under the CC-BY 4.0 license. Most of the values are missing, with only 10% present. Sequences are relatively short, ranging from 8 to 70 elements, with a median length of 38. The target variable indicates whether the patient developed sepsis, meaning it’s binary and highly unbalanced, so we measure AUROC. This dataset is also used by the original NCDE paper kidger2020neural, making it a valuable addition to our benchmark. Following the NCDE paper, we only consider the first 72 hours of a patient’s stay.
  • Two datasets from the Time Series Classification archive  [^5] https://www.timeseriesclassification.com/ middlehurst2024bake (with no explicit license specified): the U-Wave Gesture Library liu2009uwave and the Insect Sound dataset [^6] https://www.timeseriesclassification.com/description.php?Dataset=InsectSound. These datasets were processed identically: the timestamps were taken to be evenly spaced between 0 and $1$, and 30% of each sequence was replaced with NaNs. They both have a multiclass label, so we use the Accuracy metric.
  • A synthetic Pendulum dataset, with the dampening coefficient as the target osin2024ebes (with no explicit license specified) — regression task. Sequences are irregular, containing 10% missing values, with lengths varying from approximately 200 to 400 elements, the median length being 315.
@@ C.2 Backbones
The recurrent networks are all single-layer and one-directional, with 32 hidden channels. For GRU chung2014empirical we use the standard PyTorch implementation:
@@
We use the Pytorch implementation of the Transformer Encoder layer. The two versions differ in their positional embeddings: RoFormer uses rotary embeddings su2021roformer (provided by TorchTune), and TempFormer uses temporal sine-based embeddings, as proposed by the Transformer Hawkes paper zuo2020transformer.
@@
The Mamba model we use is Mamba2 gu2023mamba from the Mamba SSM library  [^7] https://github.com/state-spaces/mamba, which we employ without significant modifications.
@@
To reproduce the results as closely as possible, we use TorchCDE  [^8] https://github.com/patrick-kidger/torchcde for the Neural CDE method kidger2020neural. Specifically, we implemented the example for irregular sequence classification from their repository. For Neural RDE morrill2021neural, we pre-process the data using the original Signatory package, provided by the authors  [^9] https://github.com/patrick-kidger/signatory, with depth set to 2.
@@
As the model for GRU-ODE de2019gru, we simply take a Neural CDE with the Sync-NF vector field, without time scaling.
@@
DeNOTS is slower than other non-ODE methods; however, we argue this is an implementation issue. Neural ODEs received much less attention than State-Space Models, Transformers, or RNNs. We use the TorchODE library lienen2022torchode. Although it is faster than the original implementation chen2018neural, it still requires work, being built almost single-handedly by its main contributor.
@@ C.3 Pipeline
We provide our repository with automated scripts to download and preprocess all datasets, as well as to train and evaluate all the considered models, using popular frameworks such as Pytorch Ansel_PyTorch_2_Faster_2024 (Caffe2 license), Pytorch Lightning Falcon_PyTorch_Lightning_2019 (Apache License 2.0) and Hydra Yadan2019Hydra (MIT License) to make the process familiar to most AI researchers. The YAML configs, containing all the hyperparameters, are also provided in the repository to facilitate reproducibility. The documentation clearly explains the steps necessary to reproduce all considered experiments.
@@ C.3.1 Training
All training uses the Adam method kingma2014adam, with the learning rate fixed to $10^{-3}$ and other parameters left default. The whole model, consisting of the head and the backbone, is trained end-to-end. The head consists of a linear layer and an activation function. We do not set an upper bound for epochs, stopping only when the validation metric stops improving. The choice of head’s activation, loss, and specific metrics depends on the considered task, as outlined in Table 12.
@@ C.3.2 Embeddings
Before passing $\{\mathbf{x}_{k}\}_{k}$ to the backbone, we apply Batch Normalization ioffe2015batch. For a fairer baseline comparison, time intervals $t_{k}-t_{k-1}$ are also included in the embeddings in the same way as other numerical features.
@@
C.4 Full Results
In this section, we present the full results of all our models, which were used to build the ranks from Table 9. The results are given in Table 13.
Backbone UWGL InsectSound Pendulum Sepsis
GRU $0.5\pm 0.1$ $0.3\pm 0.2$ $0.73\pm 0.03$ $0.838\pm 0.004$
RoFormer $0.74\pm 0.04$ $0.29\pm 0.02$ $0.61\pm 0.02$ $0.924\pm 0.004$
Mamba $0.71\pm 0.07$ $0.41\pm 0.03$ $0.61\pm 0.03$ $0.829\pm 0.004$
GRU-ODE $0.4\pm 0.3$ $0.18\pm 0.03$ $0.6\pm 0.05$ $0.925\pm 0.003$
C.5 Expressivity discussion
Here, we explain why measuring expressivity with the downstream metric is a valid approach. Theoretically, the total error can be decomposed into three terms guhring2020expressivity:
@@ C.5 Expressivity discussion
Next, it is the consensus that the more expressive models suffer more from overfitting hawkins2004problem, i.e. generalize worse. We verify that our models adhere to this by performing the following experiment. We construct a smaller version of the Pendulum dataset; its training set is 1/32 of the original one, and compare the DeNOTS model with various values of $D$ on this benchmark. The results are provided in Table 14. The overfitting effect is observed very clearly: the ”shallow“ $D=1$ model achieves higher metrics on the test set than $D=10$, and the ranking is reversed on the train set.
@@ C.6 NFE-Correlation experiments
C.7 Negative Feedback Strength
@@ C.8.2 Sine-2 Dataset
C.9 Full attack results
@@
Additionally, we test how the value of $D$ affects the robustness of our models. This is presented in Figures $25$. Increasing depth lowers robustness, which is in line with the consensus. Prior work indicates that the more accurate models with higher metrics often perform worse under adversarial attacks than those with lower metrics su2018robustness. Increased accuracy often requires increased expressiveness and sensitivity, which directly impacts robustness. Consequently, this demonstration supports our claim that larger values of $D$ correspond to more expressive models.